OpenStudy (anonymous):

34/6= 2x+1/3 Its a propportion the slashes are fraction bars solve for x

OpenStudy (anonymous):

2x = 34/6 - 1/3 2x = (34-2)/6 2x=32/6 x=32/12=8/3

OpenStudy (anonymous):

ok if its a proporian its supposed to be a whole number not a fraction

OpenStudy (anonymous):

8/3 = 2.67

OpenStudy (anonymous):

It is an equation, solving for x; and the answer is what it is 8/3. Good job dindatc.